OpenStudy (anonymous):

Solve the following system of equations by the substitution method. 8x = 2y + 5 3x = y + 7 What is the solution set? {(-41/2, -9/2)} {(-9/2, -41/2)} Ø

OpenStudy (anonymous):

x=-9/2 y=-41/2

OpenStudy (anonymous):

how did u get that

OpenStudy (goformit100):

{(-9/2, -41/2)}

OpenStudy (anonymous):

should i tell you the procedure

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

see you can solve this by elimination method

OpenStudy (anonymous):

multiply second equation on both sides by 2

OpenStudy (anonymous):

6x=2y+14

OpenStudy (anonymous):

From first equation, find 2y ie 2y = 8x - 5 Now multiply second equation by 2 6x = 2y + 14 Now put the value of 2y in it, 6x = 8x - 5 + 14 -2x = 9 x = -9/2 and y = -41/2 is the answer...

OpenStudy (anonymous):

ook i c u to now

OpenStudy (anonymous):

got it???

OpenStudy (anonymous):

kinda not all to much tho

OpenStudy (anonymous):

use substitution substitute this y=3x-7 8x=2(3x-7)+5 8x=6x-14+5 8x=6x-9 -6x -6x _________________ 2x=-9 x= -9/2 3(-9/2)=y+7 substitute -9/2 for x -27/2= y+7 -7 -7 ______________ -41/2 =y (-9/2, -41/2)-------> coordinates and answer

OpenStudy (anonymous):

do you have it??

OpenStudy (anonymous):

yes thx alot

OpenStudy (anonymous):

in which step yu are getting problem?????

OpenStudy (anonymous):

NP!!!

OpenStudy (anonymous):

it was the switching things out

OpenStudy (anonymous):

when your do a problem like such if you see a variable on its own thats a clue to know its best to use substitution and in you problem above you have 8x = 2y + 5 3x = y + 7 Look at each problem closely, here you have a y that stands alone with out a number in front of it...so its best to get that y by itself and solve x... How do you do that? First I would subtract y as such 3x=y+7 -y -y ____________ -y+3x=7 But remember you have to get y by itself...so then I would subtract 3x -y+3x=7 -3x -3x _______________ -y=-3x+7 Now we have y by itself But its negative, BIG PROBLEM so in order to get Y positive you have to divide each side by -1 \[\frac{-y}{-1}=\frac{-3x+7}{-1}\] that would give you y=3x-7

OpenStudy (anonymous):

thx i wote it down

OpenStudy (anonymous):

NP